OpenStudy (anonymous):

2(x-6)<4 and 14+x+2≥20

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

then solve for x

OpenStudy (anonymous):

I solved them separately

OpenStudy (anonymous):

and then wrote the answer out showing both in one....

OpenStudy (anonymous):

okay can i see

OpenStudy (anonymous):

2(x-6) < 4 so multiply 2 to both x and -6 you get 2x-12 < 4

OpenStudy (anonymous):

you have to solve for x on both equations

OpenStudy (anonymous):

get x alone so add 12 to each side 2x < 16 and then divide each side by 2 you get x < 8

OpenStudy (anonymous):

the second equation so \[16+x \ge 20\]

OpenStudy (anonymous):

subtract 16 from each side \[x \ge 4\]

OpenStudy (anonymous):

so what would 7x+14-4<4

OpenStudy (anonymous):

\[4 \le x < 8\]

OpenStudy (anonymous):

can u do the one on top

OpenStudy (anonymous):

the 7x + 14 one?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

where did you get that from? is it a completely separate problem?

OpenStudy (anonymous):

it is a differnt problem but im haveing troblue with it too and ur good at them so i thought you could help

OpenStudy (anonymous):

7x+14-4<4 becomes 7x+10< 4 then subtract 10 from each side, you get 7x< -6 divide each side by 7 x< -6/7

OpenStudy (anonymous):

how would u graph this

OpenStudy (anonymous):

oy! ok dashed line at (-6/7,0) and shade down (or under it) does that sound right